# Lagrangian chaos and the enstrophy cascade in Ekman-Navier-Stokes two-dimensional turbulence

### v1

@@ I Introduction
The distinctive feature of 2D turbulence lies in the existence of two quadratic inviscid invariants, energy and enstrophy, which leads to a radically different phenomenology compared to classical three-dimensional (3D) turbulence [Kraichnan1967, Boffetta2012]. In 2D turbulence, energy flows toward large scales, producing an inverse cascade, while enstrophy is transported by a direct cascade toward the small scales. The statistical properties of the two cascades have been studied in detail both numerically [Frisch1984, Boffetta2010] and experimentally [Zhu2023, Paret1997, Rivera1998] and 2D Navier-Stokes turbulence has become a prototype for two-dimensional turbulent flows and, more in general, for other systems that exhibit a double cascade, such as wave turbulence [zhu2023direct, korotkevich2008simultaneous, nazarenko2011wave].
The interaction of the two-dimensional flow with the three-dimensional environment introduces additional terms to the 2D Navier-Stokes equation which can be often represented by a linear dissipation. This is the case of the bottom friction induced by the container in a thin layer of fluid [Boffetta2005], the air friction on flowing soap film [Rivera2000] or the Ekman friction in geophysical flows [Vallis2017]. Linear friction has also the important role of removing the energy transferred at large scales by the inverse cascade and thus to allow the system to reach a statistically stationary state [Boffetta2012].
The effect of friction on the direct cascade is more subtle, since it removes the enstrophy at all scales and thus produces a cascade with a non-constant enstrophy flux. One important consequence is the correction of the spectral slope in the cascade which becomes steeper than the dimensional prediction. Moreover, in these conditions the vorticity at small-scale can be considered passive and one can exploit the results obtained for the statistics of a passive scalar with finite lifetime transported by a smooth flow [Chertkov1998, Nam1999]. On this basis, one expects that the vorticity statistics can be expressed in terms of the friction coefficient and the Lagrangian statistics of stretching rates commonly characterized in terms of Finite Time Lyapunov Exponents (FTLE) [Nam2000, Boffetta2002].
@@ II.1 Phenomenology of 2D turbulence with friction
We assume that the forcing is localized around a characteristic injection scale $\ell_{I}=2\pi\sqrt{\varepsilon_{I}/\eta_{I}}$, while viscosity and friction dissipation define a small viscous scale $\ell_{\nu}=2\pi\sqrt{\varepsilon_{\nu}/\eta_{\nu}}$ and a large friction scale $\ell_{\alpha}=2\pi\sqrt{\varepsilon_{\alpha}/\eta_{\alpha}}$. A scale separation, $\ell_{\nu}\ll\ell_{I}\ll\ell_{\alpha}$, is achieved in the limit of small viscosity and friction coefficient ($\nu\to 0$, $\alpha\to 0$). In these conditions, the dynamics of 2D turbulent flows display the development of a direct enstrophy cascade, characterized by a constant flux of enstrophy along the scales $\ell_{\nu}\ll\ell\ll\ell_{I}$ and an inverse energy cascade with a constant energy flux in the scales $\ell_{I}\ll\ell\ll\ell_{\alpha}$ [Kraichnan1967, Boffetta2012].
The assumption of a constant enstrophy flux in the range of scales of the direct cascade leads to the prediction of a power law energy spectrum $E(k)\sim k^{-3}$ with possible logarithmic corrections [Kraichnan1967, Kraichnan1971]. However, numerical and experimental studies [Nam2000, Bernard2000, Boffetta2002, Boffetta2005, Valadao2025] have shown that the presence of friction causes the steepening of the energy spectrum $E(k)\sim k^{-(3+\xi)}$, with a friction-dependent correction to the spectral slope $\xi(\alpha)>0$. In these conditions, vorticity fluctuations at small scales can be considered as being transported like a passive scalar by a smooth, large-scale, chaotic velocity field. The combined effect of chaotic advection and linear damping causes anomalous scaling and small-scale intermittency in the statistics of the vorticity field [Nam2000, Bernard2000].
The phenomenon of the steepening of the spectrum can be interpreted as a result of the dissipation of enstrophy by friction, which causes a reduction of the enstrophy flux in the direct cascade. In the limit of strong friction ($1/\alpha\to 0$), an estimate of the residual dissipation of enstrophy due to viscous forces at small scales can be obtained by means of a classic argument by Fjørtoft[Fjortoft1953]. Using the dimensional estimate for the palinstrophy $P\sim Z/\ell^{2}_{\nu}$, the viscous dissipation rate of enstrophy can be written as follows:
@@ II.2 Finite Time Lyapunov Exponents and spectral correction
According to the analogy between the dynamics of vorticity fluctuations at small scales in the presence of friction and the chaotic advection of a passive scalar field with finite lifetime proposed in [Nam2000, Bernard2000], the correction $\xi$ to the slope of the energy spectrum in the direct cascade is related to the statistics of FTLE in terms of the Cramér function as
@@ II.3 Predictions for the Lyapunov exponent
In the high-friction regime, it is possible to derive a prediction for the Lyapunov exponent and for the statistics of the FTLE in terms of the friction coefficient $\alpha$ and the enstrophy injection rate $\eta_{I}$. In this limit, the advection and viscous terms in the Ekman-Navier-Stokes equation (1) become negligible [valadao2025high] and the dynamics of the large-scale vorticity is dominated by friction and forcing terms. In the case of Gaussian random forcing, the dynamics can be traced back to an Ornstein-Uhlenbeck process with a correlation time given by the inverse of the friction coefficient. The resulting vorticity field is a short-correlated in time, Gaussian field with a temporal correlation function given by
@@
In the limit $1/\alpha\to 0$ one recovers the Kraichnan ensemble [kraichnan1968small], i.e., a white-in-time, Gaussian field with temporal correlation function given by [valadao2025high]
@@
where $D_{1}$ is a constant related to two-particle dispersion [falkovich2001particles]. By matching the time integral of (18) with that of (19), one gets the relations $D_{1}=Z/4\alpha=\eta_{I}/8\alpha^{2}=Z^{2}/2\eta_{I}$, where the last two equalities follow from (8).
@@
where $\lambda=2D_{1}$ [falkovich2001particles]. This leads to the prediction for the Lyapunov exponent in the high-friction asymptotics of the Ekman-Navier-Stokes system:
@@ III Numerical simulations and results
Direct numerical simulations of Eq. (1) have been performed by means of a pseudo-spectral code with a fourth-order Runge–Kutta time scheme implemented on GPUs. A detailed description of the code is reported in [Valadao2025]. The computational domain is a square of size $L_{x}=L_{y}=2\pi$ with periodic boundary conditions. It is discretized with a uniform grid of resolution $N=N_{x}=N_{y}=1024$. The flow is sustained by a Gaussian random forcing with zero mean and white-in-time correlations, acting within a narrow spherical shell of thickness $\delta k$ centered at $k_{I}$ in wavenumber space. This forcing protocol injects energy and enstrophy at constant rates $\varepsilon_{I}$ and $\eta_{I}$, respectively, which are related as $\eta_{I}\approx\varepsilon_{I}k_{I}^{2}$ for $k_{I}\gg\delta k$. We performed a set of 25 different simulations, varying the friction coefficient $\alpha$ while keeping the other parameters constant. The numerical values of the simulation parameters are reported in Table 1.
Together with the Navier-Stokes equation, for each simulation, we also integrated an ensemble of $N_{L}=100$ Lagrangian trajectories starting from initial random positions. The numerical integration of the trajectories $\dot{\mathbf{x}}=\mathbf{u}(\mathbf{x},t)$ is performed via a fourth-order Runge-Kutta scheme with linear interpolation of the fluid velocity field at the position $\mathbf{x}$. The computation of the FTLE $\gamma_{T}$ is performed according to standard methods [Benettin1980] by following along each Lagrangian trajectory the time evolution of an infinitesimal displacement $\delta\mathbf{x}$ which is determined by the local velocity gradients $\delta\dot{\mathbf{x}}=(\nabla\mathbf{u})^{\intercal}\delta\mathbf{x}$. The $\gamma_{T}$ are computed according to Eq. (9) and their statistics are collected for different time intervals $T$. The Lyapunov exponent is obtained as the average $\left\langle\gamma_{T}\right\rangle=\lambda$. To quantify the effects of the finite resolution on the computation of the spectral correction, we performed an additional set of 15 simulations of Eq. (1) ($\alpha\in[0.16,2.56]$) at the increased resolution $N=8192$. All simulation parameters are kept constant except from the viscosity, which is reduced to $\nu=5\times 10^{-6}$.
@@ III.2 Spectral correction
In Figure 5, we show the energy spectra for the different values of the friction coefficient. As expected, we observe a power-law scaling at intermediate wavenumbers where $E(k)\sim k^{-(3+\xi)}$. To determine precisely the spectral correction, we measured the exponent $\xi$ from the power-law scaling of the enstrophy flux $\Pi_{Z}(k)\sim k^{-\xi}$ using the standard logarithmic derivative method. As discussed in detail in [Valadao2025], this procedure gives a more robust evaluation of $\xi$, in particular in the case of small friction. Nonetheless, we found that the simulations at resolution $N=1024$ do not allow a precise determination of the exponent $\xi$ since its value is affected by a relatively large uncertainty which depends on the range of wavenumbers on which it is measured. To have a more accurate measure of $\xi$, we therefore used the simulations at resolution $N=8192$.
@@ IV Conclusions
Our findings reveal that, in the presence of strong linear friction, the statistics of the FTLE in two-dimensional turbulence are mostly determined by the dynamics at the forcing scale. This suggests the possibility that the statistics could depend non-trivially on the details of the external forcing. It would be interesting to systematically investigate different forcing protocols in future studies.

### v2

@@ I Introduction
The distinctive feature of 2D turbulence lies in the existence of two quadratic inviscid invariants, energy and enstrophy, which leads to a radically different phenomenology compared to classical three-dimensional (3D) turbulence [1, 2]. In 2D turbulence, energy flows toward large scales, producing an inverse cascade, while enstrophy is transported by a direct cascade toward the small scales. The statistical properties of the two cascades have been studied in detail both numerically [3, 4] and experimentally [5, 6, 7] and 2D Navier-Stokes turbulence has become a prototype for two-dimensional turbulent flows and, more in general, for other systems that exhibit a double cascade, such as wave turbulence [8, 9, 10].
The interaction of the two-dimensional flow with the three-dimensional environment introduces additional terms to the 2D Navier-Stokes equation which can be often represented by a linear dissipation. This is the case of the bottom friction induced by the container in a thin layer of fluid [11], the air friction on flowing soap film [12] or the Ekman friction in geophysical flows [13]. Linear friction has also the important role of removing the energy transferred at large scales by the inverse cascade and thus to allow the system to reach a statistically stationary state [2].
The effect of friction on the direct cascade is more subtle, since it removes the enstrophy at all scales and thus produces a cascade with a non-constant enstrophy flux. One important consequence is the correction of the spectral slope in the cascade which becomes steeper than the dimensional prediction. Moreover, in these conditions the vorticity at small-scale can be considered passive and one can exploit the results obtained for the statistics of a passive scalar with finite lifetime transported by a smooth flow [14, 15]. On this basis, one expects that the vorticity statistics can be expressed in terms of the friction coefficient and the Lagrangian statistics of stretching rates commonly characterized in terms of Finite Time Lyapunov Exponents (FTLE) [16, 17].
@@ II.1 Phenomenology of 2D turbulence with friction
We assume that the forcing is localized around a characteristic injection scale $\ell_{I}=2\pi\sqrt{\varepsilon_{I}/\eta_{I}}$, while viscosity and friction dissipation define a small viscous scale $\ell_{\nu}=2\pi\sqrt{\varepsilon_{\nu}/\eta_{\nu}}$ and a large friction scale $\ell_{\alpha}=2\pi\sqrt{\varepsilon_{\alpha}/\eta_{\alpha}}$. A scale separation, $\ell_{\nu}\ll\ell_{I}\ll\ell_{\alpha}$, is achieved in the limit of small viscosity and friction coefficient ($\nu\to 0$, $\alpha\to 0$). In these conditions, the dynamics of 2D turbulent flows display the development of a direct enstrophy cascade, characterized by a constant flux of enstrophy along the scales $\ell_{\nu}\ll\ell\ll\ell_{I}$ and an inverse energy cascade with a constant energy flux in the scales $\ell_{I}\ll\ell\ll\ell_{\alpha}$ [1, 2].
The assumption of a constant enstrophy flux in the range of scales of the direct cascade leads to the prediction of a power law energy spectrum $E(k)\sim k^{-3}$ with possible logarithmic corrections [1, 18]. However, numerical and experimental studies [16, 19, 17, 11, 20] have shown that the presence of friction causes the steepening of the energy spectrum $E(k)\sim k^{-(3+\xi)}$, with a friction-dependent correction to the spectral slope $\xi(\alpha)>0$. In these conditions, vorticity fluctuations at small scales can be considered as being transported like a passive scalar by a smooth, large-scale, chaotic velocity field. The combined effect of chaotic advection and linear damping causes anomalous scaling and small-scale intermittency in the statistics of the vorticity field [16, 19].
The phenomenon of the steepening of the spectrum can be interpreted as a result of the dissipation of enstrophy by friction, which causes a reduction of the enstrophy flux in the direct cascade. In the limit of strong friction ($1/\alpha\to 0$), an estimate of the residual dissipation of enstrophy due to viscous forces at small scales can be obtained by means of a classic argument by Fjørtoft[21]. Using the dimensional estimate for the palinstrophy $P\sim Z/\ell^{2}_{\nu}$, the viscous dissipation rate of enstrophy can be written as follows:
@@ II.2 Finite Time Lyapunov Exponents and spectral correction
According to the analogy between the dynamics of vorticity fluctuations at small scales in the presence of friction and the chaotic advection of a passive scalar field with finite lifetime proposed in [16, 19], the correction $\xi$ to the slope of the energy spectrum in the direct cascade is related to the statistics of FTLE in terms of the Cramér function as
@@ II.3 Predictions for the Lyapunov exponent
In the high-friction regime, it is possible to derive a prediction for the Lyapunov exponent and for the statistics of the FTLE in terms of the friction coefficient $\alpha$ and the enstrophy injection rate $\eta_{I}$. In this limit, the advection and viscous terms in the Ekman-Navier-Stokes equation (1) become negligible [22] and the dynamics of the large-scale vorticity is dominated by friction and forcing terms. In the case of Gaussian random forcing, the dynamics can be traced back to an Ornstein-Uhlenbeck process with a correlation time given by the inverse of the friction coefficient. The resulting vorticity field is a short-correlated in time, Gaussian field with a temporal correlation function given by
@@
In the limit $1/\alpha\to 0$ one recovers the Kraichnan ensemble [23], i.e., a white-in-time, Gaussian field with temporal correlation function given by [22]
@@
where $D_{1}$ is a constant related to two-particle dispersion [24]. By matching the time integral of (18) with that of (19), one gets the relations $D_{1}=Z/4\alpha=\eta_{I}/8\alpha^{2}=Z^{2}/2\eta_{I}$, where the last two equalities follow from (8).
@@
where $\lambda=2D_{1}$ [24]. This leads to the prediction for the Lyapunov exponent in the high-friction asymptotics of the Ekman-Navier-Stokes system:
@@ III Numerical simulations and results
Direct numerical simulations of Eq. (1) have been performed by means of a pseudo-spectral code with a fourth-order Runge–Kutta time scheme implemented on GPUs. A detailed description of the code is reported in [20]. The computational domain is a square of size $L_{x}=L_{y}=2\pi$ with periodic boundary conditions. It is discretized with a uniform grid of resolution $N=N_{x}=N_{y}=1024$. The flow is sustained by a Gaussian random forcing with zero mean and white-in-time correlations, acting within a narrow spherical shell of thickness $\delta k$ centered at $k_{I}$ in wavenumber space. This forcing protocol injects energy and enstrophy at constant rates $\varepsilon_{I}$ and $\eta_{I}$, respectively, which are related as $\eta_{I}\approx\varepsilon_{I}k_{I}^{2}$ for $k_{I}\gg\delta k$. We performed a set of 25 different simulations, varying the friction coefficient $\alpha$ while keeping the other parameters constant. The numerical values of the simulation parameters are reported in Table 1.
Together with the Navier-Stokes equation, for each simulation, we also integrated an ensemble of $N_{L}=100$ Lagrangian trajectories starting from initial random positions. The numerical integration of the trajectories $\dot{\mathbf{x}}=\mathbf{u}(\mathbf{x},t)$ is performed via a fourth-order Runge-Kutta scheme with linear interpolation of the fluid velocity field at the position $\mathbf{x}$. The computation of the FTLE $\gamma_{T}$ is performed according to standard methods [25] by following along each Lagrangian trajectory the time evolution of an infinitesimal displacement $\delta\mathbf{x}$ which is determined by the local velocity gradients $\delta\dot{\mathbf{x}}=(\nabla\mathbf{u})^{\intercal}\delta\mathbf{x}$. The $\gamma_{T}$ are computed according to Eq. (9) and their statistics are collected for different time intervals $T$. The Lyapunov exponent is obtained as the average $\left\langle\gamma_{T}\right\rangle=\lambda$. To quantify the effects of the finite resolution on the computation of the spectral correction, we performed an additional set of 15 simulations of Eq. (1) ($\alpha\in[0.16,2.56]$) at the increased resolution $N=8192$. All simulation parameters are kept constant except from the viscosity, which is reduced to $\nu=5\times 10^{-6}$.
@@ III.2 Spectral correction
In Figure 5, we show the energy spectra for the different values of the friction coefficient. As expected, we observe a power-law scaling at intermediate wavenumbers where $E(k)\sim k^{-(3+\xi)}$. To determine precisely the spectral correction, we measured the exponent $\xi$ from the power-law scaling of the enstrophy flux $\Pi_{Z}(k)\sim k^{-\xi}$ using the standard logarithmic derivative method. As discussed in detail in [20], this procedure gives a more robust evaluation of $\xi$, in particular in the case of small friction. Nonetheless, we found that the simulations at resolution $N=1024$ do not allow a precise determination of the exponent $\xi$ since its value is affected by a relatively large uncertainty which depends on the range of wavenumbers on which it is measured. To have a more accurate measure of $\xi$, we therefore used the simulations at resolution $N=8192$.
@@ IV Conclusions
Our findings reveal that, in the presence of strong linear friction, the statistics of the FTLE in two-dimensional turbulence are mostly determined by the dynamics at the forcing scale. This suggests the possibility that the statistics could depend non-trivially on the details of the external forcing. It would be interesting to systematically investigate different forcing protocols in future studies. On a different note, it would also be interesting to apply the same approach to different systems. The hypotheses of the theoretical framework used here seem quite general, so in principle they should be applicable to other equations. However, we are not aware of any other physical system displaying both the necessary scaling properties and the correct dynamics. Magneto-hydrodynamic turbulence for example, has an advective dynamics (which is necessary for the connection with the Lagrangian FTLE) but it does not exibit a smooth direct cascade regime [26]. Wave turbulence models [10], on the other hand, are not in general dominated by an advective non-linear term driving the downscale transfer of the cascading quantity. Although we cannot exclude that similar approaches can be used for turbulent wave equations or closely related quantum turbulence models (e.g., the Gross-Pitaevskii model [8]), the Lagrangian side of the theoretical argument seems hard to justify.